OpenStudy (caitlinnr14):

HELP WILL GIVE MEDAL FOR BEST ANSWER!!!!!! Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7. y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (anonymous):

let y = x + k putting x = 6 and y = -2 -2=6+k k = -8 answer is y = x -8

OpenStudy (caitlinnr14):

thx